Draw a consumer budget constraint and indifference curves for Pepsi and Pizza. Find out

the optimal consumption choice when consumer has income of $2000 and the price of Pepsi is $5 per bottle and the price of per pizza is $10.What is the marginal rate of substitution at this optimum?


1
Expert's answer
2022-03-29T12:13:54-0400

The solution

The line AF in the diagram below shows the various combinations of commodities that a consumer can acquire. This is the budget line or budget limit. Given the prices of the goods and the consumer's income, it depicts the various combinations of Pizza and Pepsi that a consumer could purchase. Points A, B, C, D, E, and F represent these combinations.


The indifference curve associated with the budget line above is as shown below.


"Budget=P_1 Q1_\u200b + P_2\u200bQ_2\u200bP \n1\n\u200b\n \n\u200b\n \u200b"

"2000=5x+10y"

"\u2223MRS \nxy\n\u200b\n \u2223= \n\u03b4x\n\u03b4y\n\u200b\n = \n\\frac{MUx}{MUy}\n\u200b\n \n\u200b"

"= \n\\frac{MUx}{MUy}\n\u200b\n \n\u200b\n = \n\\frac{-5}{-10}\n\u200b\n =0.5"